OpenStudy (anonymous):

A test of the prototype of a new automobile shows that the minimum distance for a controlled stop from 94 km/h to zero is 41 m. Find the acceleration, assuming it to be constant as a fraction of the free-fall acceleration. The acceleration due to gravity is 9.81 m/s 2 . Answer in units of g How much time does the car take to stop? Answer in units of s

OpenStudy (anonymous):

@Agent47 can you help me right here

OpenStudy (agent47):

d=41m V_i=94km/h = 94,000m/h = 94000m/60s = 1,566.67 m/s V_f=0 a-?

OpenStudy (agent47):

a=c*g d=V_i*t+(1/2)(c*g)(t^2) a=(V_f-V_i)/t 41=1,566.67+(c*g*t^2)/2 c*g=(0-1,566.67)/t

OpenStudy (agent47):

-3051.34 = c*g*t^2 -1566.67 = c*g*t

OpenStudy (agent47):

dividing the two equations gives: -3051.34/-1566.67=cgt^2/(cgt) t = 1.95

OpenStudy (agent47):

uhmm c*g = -1,566.67/t c*g = -1,566.67/1.95 a = c*g = -803.42

OpenStudy (agent47):

t=1.95s acceleration is: a = -81.98*g

OpenStudy (agent47):

I got -81.98 from dividing -803.42 by 9.81

OpenStudy (anonymous):

okay im writing it all down

OpenStudy (anonymous):

this is how its done